OpenStudy (anonymous):

I'm having trouble changing limit \[\int_{0}^{1} x(\sqrt(x^2 +4))\] my plan is to solve the integral using \[x=2tan\theta\] \[dx=2sec^{2}\theta d\theta\]

OpenStudy (anonymous):

0=2tan theta so theta=0 1=2sin theta

OpenStudy (anonymous):

no it's 2=sin theta

OpenStudy (anonymous):

put x^2+4 = t 2*x*dx = dt

OpenStudy (anonymous):

what about the limit for some reason I don't seem to know how to change 2=sin theta into theta=? in terms of pi...and yes i tried wolf

OpenStudy (anonymous):

\[-1<\sin \theta <1\] it cann't be 2

OpenStudy (anonymous):

well I guess I'm a having a little trouble with the integral \[16\int tan\theta sec^2 \theta sec^2 theta d\theta\]

OpenStudy (anonymous):

16\int tan\theta sec^2 \theta sec^2 \theta d\theta

OpenStudy (anonymous):

\[16\int tan\theta sec^2 \theta sec^2 \theta d\theta\]

OpenStudy (anonymous):

I think i got it u=sec^2 du= 2 tan sec^2

OpenStudy (anonymous):

\[8\int udu\]

OpenStudy (anonymous):

but still changing the limits are an issue...

OpenStudy (anonymous):

\[4 sec^4 \theta\]

OpenStudy (anonymous):

sometimes putting wrong substitution can make the question worse

OpenStudy (anonymous):

what u did last? i think its not correct .check it again

OpenStudy (experimentx):

let 2tan \theta = x , x=0, \theta = 0, for x=1 \theta = arctan(1/2) and i am getting \int 8 tan \theta \sec^3 \theta d\theta as integration

OpenStudy (anonymous):

\[2tan \theta = x , x=0, \theta = 0, for x=1 \theta = arctan(1/2)\] \[\int 8 tan \theta \sec^3 \theta d\theta\] ?

OpenStudy (anonymous):

how about dx?

OpenStudy (experimentx):

to solve this let \sec \theta = u, du = \sec \theta \tan \theta, then you have \int u^2 du = 1/3 u^3

OpenStudy (anonymous):

so the upper limit of the new limit is?

OpenStudy (zarkon):

why would you want to use trig sub for this problem?

OpenStudy (anonymous):

it's the section that I'm working on in the chapter....I'm forced to do it

OpenStudy (zarkon):

just because the problem is in that section doesn't mean that you are supposed to use that technique

OpenStudy (anonymous):

just to double check...what do you get when you do it with a different technique?

OpenStudy (zarkon):

sometimes the problems are there to test if you can see that an older technique already learned in a previous section will work better

OpenStudy (zarkon):

I would do \[u=x^2+4\]

OpenStudy (anonymous):

so yeah that's way easier

OpenStudy (anonymous):

the trig subs still bother me though

OpenStudy (anonymous):

my answer is way off...so how did you get away with not having dx= 2 sec^2 \theta d\theta?

OpenStudy (anonymous):

final answer i have \[\frac 1 3 (5\sqrt 5 +8)\]

OpenStudy (experimentx):

\[ \int x \sqrt{x^ 2 + 4} dx = \frac 1 2 \frac{(x^2 + 4)^{3/2}}{3/2}\] plugin the limits.

OpenStudy (anonymous):

\[\frac 1 2 \frac{(x^2 + 4)^{3/2}}{3/2}\] \[\frac 1 2 \frac{(1 + 4)^{3/2}}{3/2} =5\sqrt5 *\frac 2 6\] \[\frac 1 2 \frac{( 4)^{3/2}}{3/2} =8*\frac 1 3 \]

OpenStudy (anonymous):

lolz...i'm supposed to subtract them...sorry

OpenStudy (anonymous):

\[\frac5 3 \sqrt5 -\frac 8 3\] good enough i guess

OpenStudy (experimentx):

well you got \[ \frac 13 (5 \sqrt 5 - 8) \]

OpenStudy (anonymous):

or the other way around

OpenStudy (anonymous):

no that's right

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

is there something I'm missing?